OpenStudy (anonymous):

12a^3b^2+ 4a^4b^5 + 16ab^2/4ab^2 = Divide and check. Can someone tell me if I have done this correctly? 12a^3b^2 + 4a^4b^5 + 16ab^2/4ab^2 = 3a2b + a^3b^3 + 4ab ????

hartnn (hartnn):

nopes,try again 1st and 3rd term are incorrect....

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

hmm... 12a^3b^2/4ab^2 + 4a^4b^5/4ab^2 + 16ab^2/4ab^2 = 3a^2b +a^3b^3 + 4ab ( i'm still getting this number.what am i doing incorrectly?)

hartnn (hartnn):

b^2/b^2=1 and not b

OpenStudy (anonymous):

so it would be 3a^2(1) + a^3b^3 + 4(1)(1) = 3a^2(1) + a^3b^3 + 4(1)= 3a^2 + a^3b^3 + 4 ?

hartnn (hartnn):

yup,u got it correct now :)

OpenStudy (anonymous):

ty for your help :)

hartnn (hartnn):

welcome :)